OpenStudy (anonymous):

square root of 512k^2. show work

OpenStudy (anonymous):

16k * SQRT(2)

OpenStudy (anonymous):

can u explain?

OpenStudy (anonymous):

512 = 256 * 2 Square root of 512 = square root of (256 * 2) = square root (256) * square root (2) = 16 * square root of (2). square root (k^2) = k.

OpenStudy (anonymous):

im kida cofused can u use the real symbols. sorry

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

Draw function does not work right on my MAC.

OpenStudy (anonymous):

OK. I switched to my windows machine. I can draw now.

OpenStudy (anonymous):